# Global 3D Reconstruction of Clouds & Tropical Cyclones

### v1

@@ 5 Conclusion
This work has been enabled by Frontier Development Lab Earth Systems Lab (https://eslab.ai/)—a public / private partnership between the European Space Agency (ESA), Trillium Technologies, the University of Oxford and leaders in commercial AI supported by Google Cloud, Scan Computers, Nvidia Corporation and Pasteur Labs. The authors would also like to thank the reviewers and experts that provide advice throughout this project, including Julien Boussard, Gherardo Varando, Homer Durand, Milton Gomez, Johanna Mayer, Luis Gómez-Chova, Matteo Salvador, Alistair Francis, Mikolaj Czerkawski, Jacqueline Campbell, Emmanuel Johnson, Howard Barker, Sarah Brüning, Arthur Avenas, and Dominique Brunet.
@@
GOES-CloudSat 2018 - 2024 2 km @ SSP 16 31,046 72 GB
MSG-CloudSat 2004 - 2025 3 km @ SSP 11 181,653 620 GB
HIMAWARI-CloudSat 2015 - 2022 2 km @ SSP 16 57,373 181 GB

### v3

@@ 5 Conclusion
This work has been enabled by Frontier Development Lab Earth Systems Lab (https://eslab.ai/)—a public / private partnership between the European Space Agency (ESA), Trillium Technologies, the University of Oxford and leaders in commercial AI supported by Google Cloud, Scan Computers, Nvidia Corporation and Pasteur Labs. The authors would also like to thank the reviewers and experts that provide advice throughout this project, including Julien Boussard, Gherardo Varando, Homer Durand, Milton Gomez, Johanna Mayer, Luis Gómez-Chova, Matteo Salvador, Alistair Francis, Mikolaj Czerkawski, Jacqueline Campbell, Emmanuel Johnson, Howard Barker, Sarah Brüning, Arthur Avenas, Dominique Brunet, Loredana Spezzi and Alessio Bozzo.
@@
GOES-CloudSat 2018 - 2020 2 km @ SSP 16 31,046 72 GB
MSG-CloudSat 2006 - 2020 3 km @ SSP 11 181,653 620 GB
HIMAWARI-CloudSat 2015 - 2020 2 km @ SSP 16 57,373 181 GB